### v1

# The Powell Conjecture for the genus-three Heegaard splitting of the $3$-sphere

Abstract.
The Powell Conjecture states that four specific elements suffice to generate the Goeritz group of the Heegaard splitting of the $3$-sphere. We present an alternative proof of the Powell Conjecture when the genus of the splitting is $3$, and suggest a strategy for the case of higher genera.
Key words and phrases:
reducing sphere, primitive disk, the Powell Conjecture
2020 Mathematics Subject Classification:
Primary: 57K30
The first-named author is supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (NRF-2021R1F1A1060603).
The second-named author is supported by JSPS KAKENHI Grant Numbers JP20K03588, JP21H00978, and JP23H05437
The third-named author is supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (RS-2023-00275419).
1. Introduction
Let $(V,W;\Sigma)$ be a Heegaard splitting of a closed orientable 3-manifold $M$. That is, $\Sigma$ is a closed orientable surface in $M$ that decomposes $M$ into two handlebodies $V$ and $W$. The surface $\Sigma$ is called the Heegaard surface, and the genus of $\Sigma$ the genus of the splitting. The Goeritz group of the splitting is then defined to be the automorphism group of the splitting, the group of isotopy classes of orientation-preserving self-homeomorphisms of $M$ preserving each of $V$ and $W$ setwise.
Due to a result of Waldhausen [15], the Heegaard splitting of the $3$-sphere $S^{3}$ is uniquely determined by its genus up to isotopy. Thus we denote simply by $\mathcal{G}_{g}$ the Goeritz group of the genus-$g$ Heegaard splitting of $S^{3}$. It is easy to see that $\mathcal{G}_{0}=1$ and $\mathcal{G}_{1}=\mathbb{Z}/2\mathbb{Z}$, however, the difficulty level suddenly increases when the genus becomes greater than $1$. In 1933, Goeritz [11] gave a finite generating set of $\mathcal{G}_{2}$, that is a reason why the automorphism groups of Heegaard splittings are called Goeritz groups today. There has been no progress in the study of the Goeritz groups $\mathcal{G}_{g}$ around in the half century since his work. In 1980, Powell published a paper [12] claiming to have a finite generating set of $\mathcal{G}_{g}$ for $g\geq 3$, but in 2004, Scharlemann [13] pointed out that the proof contained a fatal gap. The generating set proposed by Powell still remains plausible, and so the conjecture that the set does generate $\mathcal{G}_{g}$ is now called the Powell Conjecture. We note that one of the five elements in the Powell’s generating set was shown to be redundant by Scharlemann [14], and hence the Powell Conjecture asserts that the remaining four elements generate $\mathcal{G}_{g}$.
Since the work of Scharlemann [13], the study of Goeritz groups has been developed by incorporating the techniques of various fields, such as singularity theory, dynamical systems, and geometric group theory. We refer the introduction of [8] to the reader for a brief survey of the study. Although there have been various results on Goeritz groups, the Powell Conjecture still continues to have the top priority. One of the basic principles in the study related to the conjecture is to find a nice simplicial complex on which the Goeritz group acts naturally. In fact, Scharlemann [13] gave a modern proof of the Goeritz’s classical result on $\mathcal{G}_{2}$ using the reducing sphere complex. His work was further developed by Akbas, who obtained a finite presentation of $\mathcal{G}_{2}$ in [1]. Right after that, the first-named author gave the same presentation of $\mathcal{G}_{2}$ in [7] as Akbas’s by using another complex called the primitive disk complex. Unfortunately, it seems that most arguments used in the above could not be applied directly to the case of higher genera. For example, in the Heegaard splitting of genus greater than $2$, the standard disk surgery technique, which was the clue for showing the connectivity (and contractibility) of the primitive disk complex for the genus-$2$ splitting, is shown to be no longer applicable (see [9]).
Recently, Freedman-Scharlemann [10] verified the Powell Conjecture for $\mathcal{G}_{3}$ by applying a technique from singularity theory, called graphics. On the other hand, Zupan [16] proved that if the reducing sphere complex is connected, then the Powell Conjecture is correct.
In this paper, we give a short proof for the connectivity of the reducing sphere complex for the genus-$3$ Heegaard splitting of $S^{3}$, which immediately gives an alternative proof of the Powell Conjecture for $\mathcal{G}_{3}$. Our proof uses the recent work of Campisi-Torres [5], which shows that the Heegaard surface for $S^{3}$ of genus $g>1$ is topologically minimal.
Here is the definition of the reducing sphere complex. Let $(V,W;\Sigma)$ be a genus-$g$ Heegaard splitting of $S^{3}$ with $g\geq 3$. A $2$-sphere $P$ in $S^{3}$ is called a reducing sphere for $\Sigma$ if $P\cap\Sigma$ is a single circle that is essential in $\Sigma$. If $P$ is a reducing sphere for $\Sigma$, then $P\cap V$ and $P\cap W$ are essential separating disks in $V$ and $W$, respectively. The reducing sphere complex $\mathcal{R}(\Sigma)$ is then a simplicial complex defined as follows. The vertices of $\mathcal{R}(\Sigma)$ are the isotopy classes of reducing spheres for $\Sigma$, and $k+1$ distinct vertices span a $k$-simplex if there exist representative spheres of the vertices that are pairwise disjoint.
We note that the reducing sphere complex for the genus-$2$ splitting has slightly different definition, since any two non-isotopic reducing spheres must intersect in this case. In [7], its combinatorial structure is well understood, in fact, it is a $2$-dimentional complex which deformation retracts to a tree. We will not deal with the genus-2 splitting in this paper. The main result of this work is stated as follows.
Theorem 1.1.
For the genus-$3$ Heegaard splitting of the $3$-sphere, the reducing sphere complex $\mathcal{R}(\Sigma)$ is connected.
The following is a direct consequence of Theorem 1.1 together with Zupan [16].
Corollary 1.2.
The Powell Conjecture is true in the case of genus $3$.
Throughout the paper, $\operatorname{N}(X)$ will denote a regular neighborhood of $X$ for a subspace $X$ of a space, where the ambient space will always be clear from the context.
2. Primitive disks and weak reducing disks
Throughout the section, $(V,W;\Sigma)$ will be a genus-$g$ Heegaard splitting of $S^{3}$ with $g\geq 3$. An essential disk $D$ in $V$ is called a primitive disk if there is an essential disk $E$ in $W$ such that $\partial D$ intersects $\partial E$ transversely in a single point. Such a disk $E$ is called a dual disk of $D$. The disk $E$ is also a primitive disk in $W$ with a dual disk $D$ in $V$. We note that any primitive disk is necessarily non-separating. We call the pair $(D,E)$ of a primitive disk $D$ with its dual disk $E$ simply a dual pair. Two dual pairs $(D,E)$ and $(D^{\prime},E^{\prime})$ are said to be separated if $D\cup E$ and $D^{\prime}\cup E^{\prime}$ are disjoint. Two dual pairs $(D,E)$ and $(D^{\prime},E^{\prime})$ are said to be p-connected if there exists a sequence of dual pairs $(D,E)=(D_{1},E_{1}),(D_{2},E_{2}),\ldots,(D_{n},E_{n})=(D^{\prime},E^{\prime})$ such that $(D_{i},E_{i})$ and $(D_{i+1},E_{i+1})$ are separated for each $i\in\{1,2,\ldots,n-1\}$.
Lemma 2.1.
Let $(D,E)$ and $(D^{\prime},E^{\prime})$ be dual pairs. Let $P$ and $P^{\prime}$ be reducing spheres disjoint from $D\cup E$ and $D^{\prime}\cup E^{\prime}$, respectively. Suppose that $P$ and $P^{\prime}$ are in the same component of $\mathcal{R}(\Sigma)$. Then $(D,E)$ and $(D^{\prime},E^{\prime})$ are p-connected.
Proof.
By assumption, we have a sequence $P=P_{1},P_{2},\ldots,P_{n}=P^{\prime}$ of reducing spheres such that $P_{i}$ is disjoint from $P_{i+1}$ for each $i\in\{1,2,\ldots,n-1\}$. The reducing sphere $P=P_{1}$ separates $S^{3}=V\cup_{\Sigma}W$ into two $3$-balls $B_{1}=V_{1}\cup_{\Sigma_{1}}W_{1}$ and $B^{\prime}_{1}=V^{\prime}_{1}\cup_{\Sigma^{\prime}_{1}}W^{\prime}_{1}$, where $V_{1}=V\cap B_{1}$, $W_{1}=W\cap B_{1}$, $\Sigma_{1}=\Sigma\cap B_{1}$, and $V^{\prime}_{1}=V\cap B^{\prime}_{1}$, $W^{\prime}_{1}=W\cap B^{\prime}_{1}$, $\Sigma^{\prime}_{1}=\Sigma\cap B^{\prime}_{1}$. Without loss of generality, assume that $D=D_{1}$ and $E=E_{1}$ are contained in $B_{1}=V_{1}\cup_{\Sigma_{1}}W_{1}$.
Case $1$. $P_{2}$ is contained in $V_{1}\cup_{\Sigma_{1}}W_{1}$.
We can find a dual pair $(D_{2},E_{2})$ such that $D_{2}\cup E_{2}$ is contained in $V^{\prime}_{1}\cup_{\Sigma^{\prime}_{1}}W^{\prime}_{1}$. Then $D_{2}\cup E_{2}$ is disjoint from $P_{2}$, and $(D_{1},E_{1})$ and $(D_{2},E_{2})$ are separated.
Case $2$. $P_{2}$ is contained in $V^{\prime}_{1}\cup_{\Sigma^{\prime}_{1}}W^{\prime}_{1}$.
As in the case of $P_{1}$, the reducing sphere $P_{2}$ also separates $V\cup_{\Sigma}W$ into two $3$-balls, say $V_{2}\cup_{\Sigma_{2}}W_{2}$ and $V^{\prime}_{2}\cup_{\Sigma^{\prime}_{2}}W^{\prime}_{2}$, and without loss of generality, assume that $P_{1}$ is contained in $V_{2}\cup_{\Sigma_{2}}W_{2}$. We can take a dual pair $(D_{2},E_{2})$ such that $D_{2}\cup E_{2}$ is contained in $V^{\prime}_{2}\cup_{\Sigma^{\prime}_{2}}W^{\prime}_{2}$, and is disjoint from $P_{2}$. The dual pairs $(D_{1},E_{1})$ and $(D_{2},E_{2})$ are then separated.
Inductively, we take a dual pair $(D_{i},E_{i})$ such that $D_{i}\cup E_{i}$ is disjoint from $P_{i}$, and $(D_{i-1},E_{i-1})$ and $(D_{i},E_{i})$ are separated for each $i\in\{2,\ldots,n\}$. As in the cases of $P_{1}$ and $P_{2}$, the reducing sphere $P_{n}=P^{\prime}$ separates $V\cup_{\Sigma}W$ into two $3$-balls, say $V_{n}\cup_{\Sigma_{n}}W_{n}$ and $V^{\prime}_{n}\cup_{\Sigma^{\prime}_{n}}W^{\prime}_{n}$. If $D_{n}\cup E_{n}$ and $D^{\prime}\cup E^{\prime}$ are in the same side of $P_{n}$, say in $V_{n}\cup_{\Sigma_{n}}W_{n}$, then take a dual pair $(D_{n+1},E_{n+1})$ such that $D_{n+1}\cup E_{n+1}$ lies in $V^{\prime}_{n}\cup_{\Sigma^{\prime}_{n}}W^{\prime}_{n}$. Then $(D_{n},E_{n})$ and $(D^{\prime},E^{\prime})$ are p-connected via $(D_{n+1},E_{n+1})$. If $D_{n}\cup E_{n}$ and $D^{\prime}\cup E^{\prime}$ are in the opposite sides of $P_{n}$, then $(D_{n},E_{n})$ and $(D^{\prime},E^{\prime})$ are already separated. We conclude that $(D,E)$ and $(D^{\prime},E^{\prime})$ are p-connected. ∎
A reducing sphere $P$ is said to be associated to a dual pair $(D,E)$ if $P=\partial\operatorname{N}(D\cup E)$.
Lemma 2.2.
Let $E$ be a common dual disk of two primitive disks $D$ and $D^{\prime}$. Then the dual pairs $(D,E)$ and $(D^{\prime},E)$ are p-connected.
Proof.
First suppose that $D$ is disjoint from $D^{\prime}$. Consider the reducing spheres $P$ and $P^{\prime}$ associated to $(D,E)$ and $(D^{\prime},E)$, respectively. It is shown by Zupan that if two reducing spheres intersect at most in six points in $\Sigma$ then they are in the same component of $\mathcal{R}(\Sigma)$ [16, Theorem 1.3]. Since $|(P\cap\Sigma)\cap(P^{\prime}\cap\Sigma)|=4$, $P$ and $P^{\prime}$ are in the same component of $\mathcal{R}(\Sigma)$. Then by Lemma 2.1, $(D,E)$ and $(D^{\prime},E)$ are p-connected.
Next suppose that $D$ intersects $D^{\prime}$. Let $p$ and $p^{\prime}$ be the unique intersection points of $D$ and $E$, and $D^{\prime}$ and $E$, respectively. Let $\Delta$ be a subdisk of $D^{\prime}$ cut off by an outermost arc $\delta$ of $D\cap D^{\prime}$ in $D^{\prime}$ such that $\Delta$ does not contain $p^{\prime}$. The arc $\delta$ cuts $D$ into two disks $F_{1}$ and $F_{2}$. Let $F_{1}$ be the one that contains the point $p$, and let $D_{2}=F_{1}\cup\Delta$. Since the disk $D_{2}$ intersects $E$ only in the point $p$, $(D_{2},E)$ is a dual pair. We observe that $|D_{2}\cap D^{\prime}|<|D\cap D^{\prime}|$ after a slight isotopy since at least the arc $\delta$ no longer count.
Inductively, we obtain a sequence of primitive disks $D=D_{1},D_{2},\ldots,D_{n}=D^{\prime}$ in $V$ with the common dual disk $E$ such that $D_{i}$ is disjoint from $D_{i+1}$ for each $i\in\{1,2,\ldots,n-1\}$. To show that $(D,E)$ and $(D^{\prime},E)$ are p-connected, it is enough to show that $(D_{i},E)$ and $(D_{i+1},E)$ are p-connected for each $i$, but it follows from the same argument in the first paragraph. ∎
An essential disk $D$ in $V$ is called a weak reducing disk if there is an essential disk $E$ in $W$ disjoint from $D$. The disk $E$ is also a weak reducing disk in $W$. We call $(D,E)$ (and $(E,D)$ also) a weak reducing pair. Obviously, any primitive disk is a weak reducing disk. That is, given any primitive disk $D$ in $V$ with a dual disk $E$ in $W$, the disk $\partial(\operatorname{N}(D\cup E))\cap W$ in $W$ is disjoint from $D$. A sequence $\Delta_{1}-\Delta_{2}-\cdots-\Delta_{n}$ of weak reducing disks is called a weak reducing sequence (connecting $\Delta_{1}$ to $\Delta_{n}$) if $(\Delta_{i},\Delta_{i+1})$ is a weak reducing pair for each $i\in\{1,2,\ldots,n-1\}$.
We first elaborate a short weak reducing sequence to a new one consisting only of non-separating disks in the middle.
Lemma 2.3.
Suppose that $D-E-D^{\prime}$ is a weak reducing sequence such that $E$ is a separating disk. Then there is a new weak reducing sequence $D-E^{\prime}-D^{\prime}$ or $D-E^{\prime}-D^{\prime\prime}-E^{\prime\prime}-D^{\prime}$, where $E^{\prime}$, $D^{\prime\prime}$, $E^{\prime\prime}$ are all non-separating.
Proof.
Without loss of generality, we may assume that the separating disk $E$ is contained in $W$. Let $W_{1}$ and $W_{2}$ be the two handlebodies cut off by $E$.
Case $1$. Both of the loops $\partial D$ and $\partial D^{\prime}$ lie in the same component of $\partial W-\partial E$, say both lie in $\partial W_{1}$.
Replace $E$ with a non-separating essential disk $E^{\prime}$ in $W_{2}$ to have the desired weak reducing sequence $D-E^{\prime}-D^{\prime}$.
Case $2$. The loops $\partial D$ and $\partial D^{\prime}$ lie in different components of $\partial W-\partial E$, say $\partial D$ lie in $\partial W_{1}$ while $\partial D^{\prime}$ in $\partial W_{2}$.
Compress $\Sigma$ along $D\cup E$. After the compression, let $\Sigma_{0}$ be the component containing one scar of $E$, and one scar or two scars of $D$ according to whether $D$ is separating or non-separating. Further, we maximally compress $\Sigma_{0}$ along non-separating disks to get a $2$-sphere $S$. (For the standard technique of the maximal compression, we refer [6] to the reader.) Let $\gamma$ be a circle in $S$ separating scars that came from $V$ and scars that came from $W$. The circle $\gamma$ can be regarded as a loop in $\Sigma$ and bounds essential disks in both $V$ and $W$. Thus $\gamma$ is the intersection of $\Sigma$ with a reducing sphere $P$. The reducing sphere $P$ separates $D$ and $E$, and it also separates $D$ and $D^{\prime}$ since $\partial D$ and $\partial D^{\prime}$ lie in different components of $\partial W-\partial E$.
The reducing sphere $P$ separates $S^{3}=V\cup_{\Sigma}W$ into two $3$-balls $B_{1}=V^{\prime}_{1}\cup_{\Sigma^{\prime}_{1}}W^{\prime}_{1}$ and $B_{2}=V^{\prime}_{2}\cup_{\Sigma^{\prime}_{2}}W^{\prime}_{2}$, where $V^{\prime}_{1}=V\cap B_{1}$, $W^{\prime}_{1}=W\cap B_{1}$, $\Sigma^{\prime}_{1}=\Sigma\cap B_{1}$, and $V^{\prime}_{2}=V\cap B_{2}$, $W^{\prime}_{2}=W\cap B_{2}$, $\Sigma^{\prime}_{2}=\Sigma\cap B_{2}$. Without loss of generality, we assume that $D$ is contained in $V^{\prime}_{1}$ and $D^{\prime}$ is contained in $V^{\prime}_{2}$, and further assume that the genus of $\Sigma^{\prime}_{1}$ is at least $2$. Choose a non-separating essential disk $E^{\prime}$ in $W^{\prime}_{2}$ disjoint from $P$, and then choose non-separating essential disks $D^{\prime\prime}$ in $V^{\prime}_{1}$ and $E^{\prime\prime}$ in $W^{\prime}_{1}$ such that $D^{\prime\prime}$ is disjoint from $E^{\prime\prime}$ and $D^{\prime\prime}\cup E^{\prime\prime}$ is disjoint from $P$. Then $D-E^{\prime}-D^{\prime\prime}-E^{\prime\prime}-D^{\prime}$ is the desired weak reducing sequence (see Figure 1).
Figure 1. A weak reducing sequence with $E^{\prime},D^{\prime\prime},E^{\prime\prime}$ non-separating.
A similar result holds for a weak reducing sequence $E-D-E^{\prime}$, where $D$ is a separating disk in $V$. Given any weak reducing sequence connecting $\Delta$ to $\Delta^{\prime}$, applying Lemma 2.3 repeatedly, we obtain a new weak reducing sequence connecting $\Delta$ to $\Delta^{\prime}$ in which the intermediate disks are all non-separating.
The following lemma will be used when upgrading some disks in a weak reducing sequence to primitive disks.
Lemma 2.4.
Suppose that $D-E-D^{\prime}$ is a weak reducing sequence such that $D$, $E$, and $D^{\prime}$ are all non-separating. If $D$ is disjoint from $D^{\prime}$, then there is a primitive disk $E^{\prime}$ such that $D-E^{\prime}-D^{\prime}$ is also a weak reducing sequence.
Proof.
We may assume that the disk $E$ is contained in $W$. If we compress $\Sigma$ along $E$, we get a surface $\Sigma_{0}$ containing two scars of $E$. When we further compress $\Sigma_{0}$ along $D\cup D^{\prime}$, if the two scars of $E$ lie in different components after the compression, then we have a non-separating closed surface in $S^{3}$, a contradiction. So the two scars of $E$ lie in the same component, say $\Sigma_{1}$, of the surface obtained by compressing $\Sigma_{0}$ along $D\cup D^{\prime}$. We can see that $\Sigma_{1}$ also contains either
  • the two scars of $D$ and the two scars of $D^{\prime}$, or
  • one scar of $D$ and one scar of $D^{\prime}$.
By the method of maximally compressing $\Sigma_{1}$ along non-separating disks, we get a circle $\gamma$ separating scars that came from $V$ and scars that came from $W$. Further we obtain a reducing sphere $P$ for $\Sigma$ such that $P\cap\Sigma=\gamma$ and $P$ separates $E$ and $D\cup D^{\prime}$. Let $V_{1}\cup_{\Sigma_{1}}W_{1}$ be the $3$-ball cut off by $P$ such that $E$ is contained in $W_{1}$. Replacing $E$ with a primitive disk $E^{\prime}$ in $W_{1}$, we have the desired weak reducing sequence $D-E^{\prime}-D^{\prime}$. ∎
A weak reducing sequence is called a primitive weak reducing sequence if every disk in the sequence is a primitive disk.
Lemma 2.5.
Let $D_{1}-E_{2}-D_{3}-\cdots-E_{2n}-D_{2n+1}$ be a primitive weak reducing sequence. Then there exist dual disks $\Delta_{1}$ and $\Delta_{2n+1}$ of $D_{1}$ and $D_{2n+1}$, respectively, such that the dual pairs $(D_{1},\Delta_{1})$ and $(D_{2n+1},\Delta_{2n+1})$ are p-connected.
Proof.
By compressing $\Sigma$ along $D_{1}\cup E_{2}$ and further maximally compressing the resulting surface, we can find a reducing sphere $P$ separating $D_{1}$ and $E_{2}$. Then there are dual disks $\Delta_{1}$ and $\Delta_{2}$ of $D_{1}$ and $E_{2}$, respectively, such that the dual pairs $(D_{1},\Delta_{1})$ and $(\Delta_{2},E_{2})$ are separated. Similarly there are dual disks $\Delta^{\prime}_{2}$ and $\Delta_{3}$ of $E_{2}$ and $D_{3}$, respectively, such that $(\Delta^{\prime}_{2},E_{2})$ and $(D_{3},\Delta_{3})$ are separated. The disk $E_{2}$ is a common dual disk of $\Delta_{2}$ and $\Delta^{\prime}_{2}$. By Lemma 2.2, $(\Delta_{2},E_{2})$ and $(\Delta^{\prime}_{2},E_{2})$ are p-connected. Then $(D_{1},\Delta_{1})$ and $(D_{3},\Delta_{3})$ are p-connected. See Figure 2.
Figure 2. Finding dual pairs that are p-connected.
By a similar argument, there are dual disks $\Delta^{\prime}_{3}$ and $\Delta_{5}$ of $D_{3}$ and $D_{5}$, respectively, such that $(D_{3},\Delta^{\prime}_{3})$ and $(D_{5},\Delta_{5})$ are p-connected. By Lemma 2.2 again, $(D_{3},\Delta_{3})$ and $(D_{3},\Delta^{\prime}_{3})$ are p-connected. Therefore $(D_{1},\Delta_{1})$ and $(D_{5},\Delta_{5})$ are p-connected. By an inductive argument, there exists a dual disk $\Delta_{2n+1}$ of $D_{2n+1}$ such that $(D_{1},\Delta_{1})$ and $(D_{2n+1},\Delta_{2n+1})$ are p-connected. ∎
3. Topologically minimal surfaces
Bachman defined a critical surface [2] and extended it to the notion of a topologically minimal surface [4]. Let $F$ be a surface in a $3$-manifold $M$, separating $M$ into two $3$-manifolds $V$ and $W$. The disk complex $\mathcal{D}(F)$ is a simplicial complex defined as follows. The vertices of $\mathcal{D}(F)$ are the isotopy classes of compressing disks for $F$ in $V$ and $W$, and $k+1$ distinct vertices span a $k$-simplex if they admit pairwise disjoint representatives. The surface $F$ is then called a critical surface if the set of compressing disks for $F$ in $V$ and the set of compressing disks for $F$ in $W$ can be partitioned as $\mathcal{C}_{1}~{}\dot{\cup}~{}\mathcal{C}_{2}$ and $\mathcal{C}^{\prime}_{1}~{}\dot{\cup}~{}\mathcal{C}^{\prime}_{2}$, respectively, satisfying the following properties.
  1. (1)
    There exist compressing disks $D_{i}\in\mathcal{C}_{i}$ and $E_{i}\in\mathcal{C}^{\prime}_{i}$ ($i=1,2$) such that $D_{i}\cap E_{i}=\emptyset$.
  2. (2)
    For any compressing disks $D_{i}\in\mathcal{C}_{i}$ and $E_{3-i}\in\mathcal{C}^{\prime}_{3-i}$ ($i=1,2$), $D_{i}\cap E_{3-i}\neq\emptyset$.
It is known that $F$ is critical if and only if $\pi_{1}(\mathcal{D}(F))\neq 1$. Also the following lemma is shown in [3].
Lemma 3.1.
[3, Lemma 8.5] If there exist two weak reducing disks $\Delta$ and Δsuperscriptnormal-$\Delta^{\prime}$ for $F$ such that there is no weak reducing sequence connecting $\Delta$ to Δsuperscriptnormal-$\Delta^{\prime}$, then $F$ is critical.
A surface $F$ is called a topologically minimal surface if $\mathcal{D}(F)=\emptyset$ or $\pi_{i}(\mathcal{D}(F))\neq 1$ for some $i$. The topological index of a topologically minimal surface is 00 if $\mathcal{D}(F)=\emptyset$ and the smallest $i$ such that $\pi_{i-1}(\mathcal{D}(F))\neq 1$ otherwise. A critical surface is an index $2$ topologically minimal surface. Campisi and Torres showed the following.
Theorem 3.2.
[5] The genus-$g$ Heegaard surface of $S^{3}$ is a topologically minimal surface and its topological index is $12g-$ 1.
Now we are ready to present a key lemma for the main theorem. Still we assume that $(V,W;\Sigma)$ is a genus-$g$ Heegaard splitting of $S^{3}$ with $g\geq 3$.
Lemma 3.3.
Let $D$ and $D^{\prime}$ be any non-separating weak reducing disks in $V$. Then there exists a weak reducing sequence $D=D_{1},E_{2},D_{3},$ $\ldots,E_{2n},D_{2n+1}=D^{\prime}$ such that $D_{i}$ is non-separating for each $i\in\{3,5,\ldots,2n-1\}$ and $E_{j}$ is primitive for each $j\in\{2,4,\ldots,2n\}$.
Proof.
Since $g\geq 3$, the topological index of $\Sigma$ is $2g-1\geq 5$ by Theorem 3.2. Thus $\Sigma$ is not a critical surface, and by Lemma 3.1, there exists a weak reducing sequence connecting $D$ to $D^{\prime}$. By Lemma 2.3, there is a weak reducing sequence $D=D_{1},E_{2},D_{3},\ldots,E_{2n},D_{2n+1}=D^{\prime}$ consisting of non-separating disks.
Consider a subsequence $D_{i-1}-E_{i}-D_{i+1}$. By a standard disk surgery of $D_{i-1}$ using $D_{i+1}$, we obtain a sequence of non-separating weak reducing disks $D_{i-1}=\Delta_{1},\Delta_{2},\ldots,\Delta_{m}=D_{i+1}$ such that $\Delta_{k}\cap\Delta_{k+1}=\emptyset$ for each $k\in\{1,2,\ldots,m-1\}$ and $(\Delta_{k},E_{i})$ is a weak reducing pair for each $k\in\{1,2,\ldots,m\}$. Then using Lemma 2.4, we replace $E_{i}$ with a primitive disk $\Gamma_{k}$ to obtain a weak reducing sequence $\Delta_{k}-\Gamma_{k}-\Delta_{k+1}$ (see Figure 3).
Figure 3. Upgrading $E_{i}$ to primitive disks $\Gamma_{1},\Gamma_{2},\ldots,\Gamma_{m-1}$.
By renaming and relabelling the disks, we have the desired weak reducing sequence. ∎
4. Proof of the main theorem
Throughout the section, $(V,W;\Sigma)$ will be a genus-$g$ Heegaard splitting of $S^{3}$ with $g=3$.
Lemma 4.1.
Suppose that $E-D-E^{\prime}$ is a weak reducing sequence such that $E$, $D$, and $E^{\prime}$ are all non-separating. Then $D$ is a primitive disk.
Proof.
First suppose that $E$ is disjoint from $E^{\prime}$. By compressing $\Sigma$ along $E\cup D\cup E^{\prime}$, we can find a reducing sphere $P$ for $\Sigma$ separating $E\cup E^{\prime}$ and $D$. Moreover, $D$ is contained in the side of $P$ that is the genus-$1$ summand, and hence $D$ is a primitive disk.
Next suppose that $E$ intersects $E^{\prime}$. Consider a subdisk $\Delta$ of $E^{\prime}$ cut off by an outermost arc $\delta$ of $E\cap E^{\prime}$ in $E^{\prime}$. The arc $\delta$ cuts $E$ into two disks $F_{1}$ and $F_{2}$. After a slight isotopy, both of $F_{1}\cup\Delta$ and $F_{2}\cup\Delta$ are essential disks, and at least one of them, say $F_{1}\cup\Delta$, is a non-separating disk. Obviously, $F_{1}\cup\Delta$ is disjoint from $E$ and from $D$. By compressing $\Sigma$ along $E\cup(F_{1}\cup\Delta)\cup D$, we can find a reducing sphere $P$ for $\Sigma$ separating $E\cup(F_{1}\cup\Delta)$ and $D$. Moreover, $D$ is contained in the side of $P$ that is the genus-$1$ summand, and hence $D$ is a primitive disk. ∎
Proof of Theorem 1.1.
Let $P$ and $P^{\prime}$ be any two reducing spheres. Since the genus of the splitting is $3$, there exist unique dual pairs $(D,E)$ and $(D^{\prime},E^{\prime})$ up to isotopy such that $P$ and $P^{\prime}$ are associated to $(D,E)$ and $(D^{\prime},E^{\prime})$, respectively. To show that $\mathcal{R}(\Sigma)$ is connected, that is, to find a path joining the vertices represented by $P$ and $P^{\prime}$ in $\mathcal{R}(\Sigma)$, it is enough to show that $(D,E)$ and $(D^{\prime},E^{\prime})$ are p-connected.
By Lemma 3.3 and Lemma 4.1, we have a primitive weak reducing sequence $D=D_{1},E_{2},D_{3},\ldots,E_{2n},D_{2n+1}=D^{\prime}$. By Lemma 2.5, there exist dual disks $\Delta$ and $\Delta^{\prime}$ of $D$ and $D^{\prime}$, respectively, such that $(D,\Delta)$ and $(D^{\prime},\Delta^{\prime})$ are p-connected. By Lemma 2.2, the dual pairs $(D,E)$ and $(D,\Delta)$ are p-connected, and also $(D^{\prime},\Delta^{\prime})$ and $(D^{\prime},E^{\prime})$ are p-connected. Hence $(D,E)$ and $(D^{\prime},E^{\prime})$ are p-connected. ∎
We remark that the proofs of Lemma 4.1 and Theorem 1.1 heavily depend on the fact that the genus of the splitting is $3$. But we still conjecture that the following question has an affirmative answer for the splittings of higher genera.
Question 4.2.
Given a genus-$g$ Heegaard splitting $(V,W;\Sigma)$ of the $3$-sphere with $g\geq 3$, and given any two primitive disks $D$ and $D^{\prime}$ in $V$ admitting a non-separating weak reducing disk in $W$ disjoint from $D\cup D^{\prime}$, find a primitive weak reducing sequence connecting $D$ to $D^{\prime}$.